OpenStudy (anonymous):

Find the following parts of a parabola: vertex, focus, and directrix and show work. (3x+6)^2=18y-36 18y=(3x+6)^2+36 18y=(3x+6)(3x+6)+36 18y=9x^2+36x+72 18y=9(x^2+4x+8) y=(x^2+4x+8)/2 Now what? I can't get it into parabolic form with a denominator. And, do I need to complete the square? Not sure where to go from here.

OpenStudy (anonymous):

\[y=(1/2)x^2 +2x+4\] vertex is -b/2a, f(-b/2a) =(-2,2)

OpenStudy (anonymous):

did you see your 5th line? its 9/18=1/2

OpenStudy (anonymous):

so I got y=1/2(x+2)^2+2 by dividing and completing the square... vertex (-2,2) focus (-2,5/2), and directrix y=-3/2; hopefully it's right